OpenStudy (anonymous):

Assume that y varies inversely with x and that y is 2 when x is 4. Which of the following is the equation of inverse variation?

OpenStudy (diyadiya):

Where are the equations?

OpenStudy (anonymous):

y = 2x xy = 8 y/x = 2

OpenStudy (diyadiya):

Substitute the value of x & y in the equations & find out which equation is right for example lets take the last one y/x =2 y=2 x=4 2/4=1/2≠ 2 1/2 is not equal to 2 therefore third one is wrong try the other two